OpenStudy (anonymous):

\[2x \sqrt{36-x ^{2}} \] Solve the area equation for x=3 HELP ME SOLVE!!!

OpenStudy (anonymous):

write \[2\times 3\times \sqrt{36-3^2}\] and compute

OpenStudy (anonymous):

\[6\sqrt{36+9}\]

OpenStudy (anonymous):

@satellite73 ^ ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[36-3^2=36-9=27\]

OpenStudy (anonymous):

so would it be 12\[\sqrt{5}\]

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

12\[\sqrt{5}\]

OpenStudy (anonymous):

the first step is \[2\times 3\times \sqrt{36-3^2}\]

OpenStudy (anonymous):

then you compute and get \[6\sqrt{27}\] because \(36-9=27\)

OpenStudy (anonymous):

Ok, this helps going step by step.

OpenStudy (anonymous):

since \(27=9\times 3\) you have \(\sqrt{27}=\sqrt{9}\times \sqrt{3}\) giving \[\sqrt{27}=3\sqrt{3}\]

OpenStudy (anonymous):

Would you then multiply the 3 and 6?

OpenStudy (anonymous):

so now you have \[6\times 3\sqrt{3}=18\sqrt{3}\] as a final answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Blah! Now the next part is awfully tricky.

OpenStudy (anonymous):

A(x)=2x\[\sqrt{36-x ^{2}}\]

OpenStudy (anonymous):

x=\[\sqrt{3}\]

OpenStudy (anonymous):

@satellite73 What about this?

OpenStudy (anonymous):

\[A(x)=2x\sqrt{36-x^2}\] \[A(\sqrt3)=2\sqrt{3}\times \sqrt{36-(\sqrt{3})^2}\]

OpenStudy (anonymous):

you get \[A(\sqrt{3}=2\sqrt{3}\sqrt{36-3}=2\sqrt{3}\sqrt{33}=2\sqrt{66}\]

OpenStudy (anonymous):

okay i made a mistake

OpenStudy (anonymous):

How did you end up with 66 inside the square root? I'm new to all of this. :-/

OpenStudy (anonymous):

it is \[2\sqrt{3}\sqrt{33}=2\sqrt{99}\]

OpenStudy (anonymous):

yeah i was wrong

OpenStudy (anonymous):

\[\sqrt{a}\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

How did you get the 99, just multiply 3 and 33?

OpenStudy (anonymous):

and \(3\times 33=99\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you are still not done

OpenStudy (anonymous):

Ok, let's move forward.

OpenStudy (anonymous):

since \(99=9\times 11\) you get \[\sqrt{99}=\sqrt{9}\sqrt{11}=3\sqrt{11}\]

OpenStudy (anonymous):

for a final answer of \[6\sqrt{11}\]

OpenStudy (anonymous):

This material is complicated.

OpenStudy (anonymous):

Thank you for helping. :)

OpenStudy (anonymous):

yw